An electric motor on an airplane operates on 400V volts and draws a current of 300 amperes.

(a) What is the power consumption of the motor in Watt?

(b) How much energy does the motor use during a 2 hour flight in Watt-hour?

c) If you use cells with 4V and 20Amp-hour capacity, how many cells do you need to construct the battery pack for the electric plane?

d) How many of cells must be connected in series to provide the 400V?
Answer #1

(a) Power= Voltage * current

P = VI

= 400*300

= 120000 W

(b) energy consumption = P*t

= 120000*2 = 240000 W-h

(c) power capacity of one cell = 4*20 = 80 W

number of cells required = 120000 / 80 = 1500

d) cells to be connected in series for 400 V = n

in series voltage gets added so n = 400/4 = 100